OpenStudy (anonymous):

Use the three steps to solve the problem. The length of the base of an isosceles triangle is 4 inches less than the length of one of the two equal sides of the triangles. If the perimeter is 32, find the three sides of the triangle. Place in order from smallest to largest numbers.

OpenStudy (ajprincess):

Let the length of the base be x. Then the lengths of the other two sides be will be equal to x+4 each. Perimeter=Sum f lengths of the sides 32=x+x+4+x+4 32=3x+8 Can u find x @kaiteDID?

OpenStudy (anonymous):

5

OpenStudy (ajprincess):

Hmm no. 1.Subtract 8 both sides. 2.Divide by 3 both sides. Then u will get the right answer. Can u do it?

OpenStudy (anonymous):

8

OpenStudy (ajprincess):

yup right:). Nw can u find other sides?

OpenStudy (anonymous):

12?

OpenStudy (ajprincess):

yup right:)

OpenStudy (anonymous):

you a really good at this i think i am getting a little bit better at this tank you so much

OpenStudy (anonymous):

thank*

OpenStudy (ajprincess):

yw:) Glad to knw that u r getting better in this:)